# Purification of Noisy Measurements and Faithful Distillation of Entanglement

### v1

@@ Abstract
We consider entanglement distillation in a realistic scenario with noisy operations in which quantum measurements that constitute a general quantum operation are particularly noisy. We present a protocol for purifying noisy measurements and show that with the help of the purification, imperfect local operations can be used to distill entanglement. We show that the purification protocol is robust against noise in implementation and analyze the purification in a practical realization: for measurement and gate errors up to $10\%$, we suggest that the purification with two additional qubits is cost-effective for distilling entanglement. The purification protocol is feasible with currently available quantum technologies and readily applied to entanglement applications.
Introduction. Entanglement is a resource that makes it possible to realize efficient information processing beyond the limitations of classical counterparts [1, 2]. Multipartite entangled states such as graph states can implement universal quantum computation by single-qubit measurements only [3, 4, 5]. Entanglement can establish secure communication over legitimate parties without computational assumptions [6, 7, 8, 9]. The capability of manipulating shared entangled states to useful ones by local operations of classical communication (LOCC) is a key to various entanglement applications in quantum information processing. In particular, maximally entangled states called entangled bits (ebits),
@@
are the unit of entanglement [10, 11, 12, 13, 14]. Sharing ebits allows parties to prepare other useful states, such as graph states, and to share entanglement at an arbitrary distance through quantum repeaters [15].
Distillation of entanglement is not only of practical interest to enable entanglement applications [16, 17, 18] but also of fundamental interest as it realizes an irreversible transformation from mixed or non-maximally entangled states to maximally entangled ones; note that the process is thus probabilistic [19, 20, 21, 22, 23]. Distillation of entanglement has been demonstrated experimentally [24, 25, 26, 27, 28, 29, 30, 31].
The currently available quantum technologies, the so-called noisy-intermediate-scale quantum (NISQ) technologies, contain noise in all steps, preparation, manipulation, and measurements [32]. In particular, errors are about ten times higher in a measurement than in a single-qubit gate [33, 34]. Hence, an LOCC protocol is also noisy. Note also that measurements are necessary in a construction of a general quantum operation, that can be realized by a unitary gate over a system and measurements on subsystems [35, 36].
Noisy LOCC may limit entanglement applications. For instance, the effect of noise on quantum operations has been analyzed in entanglement distillation protocols [37]. While the protocol contains two stages, bilateral controlled-NOT (CNOT) gates and measurements of single qubits, noise on gates or measurements prevents legitimate parties from distilling entanglement from weakly entangled states. The consequence reiterates the limitations of NISQ technologies towards entanglement applications such as the distribution of entangled states over a quantum network.
In this work, we present a purification protocol that probabilistically transforms a noisy measurement to a clean one, and apply it to entanglement distillation. We consider two protocols for distilling entanglement, one from unknown mixed states and the other from pure states, and show that in both cases, entanglement can be faithfully distilled with noisy measurements. The protocol is also robust against measurements and gate errors. We analyze minimal resources for practical applications of the purification protocol and show that two additional qubits suffice for distilling entanglement with noisy measurements.
@@
Purifying a noisy measurement. We begin by presenting the protocol for purifying noisy measurements. The protocol shares some similarities with advantage distillation [38, 39, 40] in key agreement protocols. Throughout, we consider a qubit measurement in the computational basis $M_{i}=|i\rangle\langle i|$ for $i=0,1$. A noisy measurement may be described by a positive-operator-valued-measure (POVM), for $i=0,1$,
@@
$\widetilde{M}_{i}=(1-p)|i\rangle\!\langle i|+p\frac{\mathds{1}}{2}$ (2)
where $p\in[0,1)$. With the decomposition $\mathds{1}=|0\rangle\langle 0|+|1\rangle\langle 1|$, one can see that $\widetilde{M}_{i}=(1-p/2)M_{i}+(p/2)M_{i\oplus 1}$, with $\oplus$ the bitwise addition, from which an error rate corresponds to $p/2$. For instance, an error rate in a NISQ system such as superconducting qubits is about $5\%$, meaning that a realistic parameter would be in a percent level, $p\approx 0.1$.
@@
Let $\rho$ denote a qubit state on which a measurement is performed. A noiseless measurement shows probabilities $\mathrm{tr}[\rho M_{i}]$, and a noisy one finds $\mathrm{tr}[\rho\widetilde{M}_{i}]$ for $i=0,1$. We here devise a protocol to obtain error-free statistics $\mathrm{tr}[\rho M_{i}]$ with noisy measurements and additional qubits without improving experimental requirements. In effect, the protocol aims to transform a noisy POVM element $\widetilde{M}_{i}$ to a clean one $M_{i}$, even with some probability.
@@
Figure 1: Detectors (red) contain noise in measurements, see Eq. (2). An error-suppressing detector (blue) that purifies a noisy measurement is devised by exploiting an interaction via a collective CNOT $V_{n}$ in Eq. (3) over $n$ qubits.
@@
Measurements on $n$ qubits are noisy. Suppose that $n$-bit outcomes are obtained, $x^{n}=(x_{1},x_{2},\cdots,x_{n})$ where $x_{k}\in\{0,1\}$ for $k=1,\cdots,n$. A POVM element describing the outcomes corresponds to
$\widetilde{M}_{x^{n}}=\widetilde{M}_{x_{1}}\otimes\cdots\otimes\widetilde{M}_{x_{n}}.$
Thus, we have the probability of outcome $x^{n}$ is found as follows,
@@
Finally, we accept measurement results only when all outcomes are identical, i.e., either $0^{n}$ or $1^{n}$; otherwise, discard. Once outcomes are accepted, a probability of outcome $i$, for $i=0,1$, is given by,
@@
which converges to a probability with a noiseless measurement, i.e., $\mathrm{tr}[\rho M_{i}]$ as $n$ tends to be large.
To see that a probability in Eq. (6) can be arbitrarily close to a result of a noiseless measurement, we rewrite the probability of outcomes $x^{n}$,
@@
$\mathrm{where}$ $\widetilde{Q}_{x_{1}}^{(n)}=\langle 0_{2}\cdots 0_{n}|V_{n}^{\dagger}\widetilde{M}_{x^{n}}V_{n}|0_{2}\cdots 0_{n}\rangle$
for $x_{1}=0,1$. For instance, one can compute the following: $V_{n}|0_{2}\cdots 0_{n}\rangle=|0^{n}\rangle\langle 0|+|1^{n}\rangle\langle 1|$, and a POVM element on the first qubit when the outcomes are identical, $x^{n}\in\{0^{n},1^{n}\}$ is given by
$\widetilde{Q}_{i}^{(n)}$ $=$ $r_{0}^{(n)}M_{i}+r_{1}^{(n)}M_{i\oplus 1}$
$\mathrm{where}$ $r_{0}^{(n)}=(1-\frac{p}{2})^{n}~{}\mathrm{and}~{}r_{1}^{(n)}=(\frac{p}{2})^{n}$ (8)
@@
for $i=0,1$. From Eq. (6), POVM elements on the first qubit when all outcomes are identical are given by,
$q^{(n)}(i)$ $=$ $\mathrm{tr}[\rho Q_{i}^{(n)}],~{}\mathrm{where}~{}$ (9)
$Q_{i}^{(n)}$ $=$ $\frac{(1-\frac{p}{2})^{n}}{p(0^{n})+p(1^{n})}M_{i}+\frac{(\frac{p}{2})^{n}}{p(0^{n})+p(1^{n})}M_{i\oplus 1}$
Since $p(0^{n})+p(1^{n})=(1-\frac{p}{2})^{n}+(\frac{p}{2})^{n}$, we have achieved a POVM element $Q_{i}^{(n)}$ which arbitrarily converges to a clean one $M_{i}$ as $n$ tends to be large. Equivalently, the fidelity between $Q_{i}^{(n)}$ and $M_{i}$ can be arbitrarily close to $1$ as $n$ increases, i.e.,
$\mathrm{tr}[Q_{i}^{(n)}M_{i}]\rightarrow 1~{}~{}\mathrm{as}~{}~{}n\rightarrow\infty$ (10)
Hence, we have shown a noisy measurement can be purified. It is also worth mentioning that the purification works for $p\in(0,1)$.
Distilling entanglement from mixed states. We now consider the distillation of entanglement with noisy local operations. We emphasize that entanglement distillation is key to various quantum information applications, such as entanglement distribution over a quantum network and measurement-based quantum computing. In what follows, we show how the purification of a noisy measurement enables existing distillation protocols to distill entanglement from all entangled states, even if local operations constructed by noisy measurements are imperfect.
@@
Let us briefly summarize the two-way distillation protocol. For copies of states shared by two parties, called Alice and Bob, the protocol begins by twirling shared and unknown states so that two parties share an isotropic state in a single-copy level, characterized as follows,
@@
where $F=\langle\phi^{+}|\rho_{F}|\phi^{+}\rangle$ is called a singlet fraction. Then, taking two copies $\rho_{F}^{A_{1}B_{1}}\otimes\rho_{F}^{A_{2}B_{2}}$, two parties apply bilateral CNOT gates, $V_{2}^{A_{1}A_{2}}$ and $V_{2}^{B_{1}B_{2}}$, see Eq. (4), and perform measurements in the second register $A_{2}B_{2}$ in the computational basis $\{M_{0},M_{1}\}$. If outcomes are equal, a state remaining in the first register $A_{1}B_{1}$ is accepted. The singlet fidelity of the accepted state is given by
@@
Figure 2: Two parties Alice and Bob apply error-suppressing detectors for distilling entanglement. For a realistic measurement error rate up to $10\%$, two additional qubits on each party immediately allow entanglement distillation from weakly entangled states up to a singlet fraction $F>0.503$.
When a measurement in the second register $A_{2}B_{2}$ is noisy, the local operations of two parties are imperfect. Consequently, entanglement cannot be distilled from weakly entangled states. For instance, for a noisy measurement with $p=0.1$, see Eq. (2), entanglement can be distilled only when $F>0.617$ [37]. No entanglement can be distilled if $p>1-1/\sqrt{2}\approx 0.29$.

### v3

@@ Abstract
We consider entanglement distillation with noisy operations in which quantum measurements that constitute a general quantum operation are particularly noisy. We present a protocol for purifying noisy measurements and show that imperfect local operations can distill entanglement. The protocol works for arbitrary noisy measurements in general and is cost-effective and resource-efficient with single additional qubit per party to resolve the distillation of entanglement. The purification protocol is feasible with currently available quantum technologies and readily applied to entanglement applications.
I Introduction
Entanglement is a resource that enables efficient information processing [1, 2] such as quantum computation [3, 4, 5] and secure communication [6, 7, 8, 9]. Manipulating shared entangled states by local operations and classical communication (LOCC) is a key to entanglement applications. In particular, an entangled bit (ebit),
@@ I Introduction
is the unit of entanglement [10, 11, 12, 13, 14].
Distilling ebits from noisy entanglement by LOCC, firstly proposed in Refs. [11, 13], makes it possible for two parties far apart to share entangled states. Once ebits are distilled, parties can extend a distance through quantum repeaters [15]. Entanglement distillation [11, 13] has been experimentally demonstrated [16, 17, 18, 19, 20, 21, 22, 23].
The results in Ref. [15, 24], however, pointed out that noisy LOCC may not be incorporated into distilling entanglement: ebits cannot be distilled from weakly entangled states. The currently available quantum technologies contain noise [25], which cannot realize entanglement distillation. Despite the importance of entanglement, no result has yet been known to enable entanglement distillation with noisy LOCC.
In this work, we present a protocol for purifying general noisy measurements and resolve the distillation of entanglement with noisy LOCC. The purification protocol incorporates the advantage distillation [26, 27] with $(n-1)$ additional qubits on two parties, respectively. We show that the purification protocol with one additional qubit, i.e., $n=2$, is cost-effective for recovering noiseless distillation. We also show that depolarization noise introduces the worst-case scenario that has the smallest range of distillability. In addition, the protocol can be extended to high dimensions. The protocol is robust against measurements and gate errors.
II Purifying Noisy Measurement
@@
We begin by presenting the protocol for purifying noisy measurements. It incorporates a cryptographic key agreement protocol called advantage distillation [26, 27, 28, 29, 30, 31]. Throughout, we consider a measurement in the computational basis $M_{i}=|i\rangle\langle i|$ for $i=0,1$. A noisy measurement may be described by a positive-operator-valued-measure (POVM),
@@
$\widetilde{M}_{i}=\mathcal{N}[|i\rangle\langle i|]~{}~{}\mathrm{such~{}that~{}~{}}\sum_{i}\mathcal{N}[|i\rangle\langle i|]=\mathds{1}.$ (2)
A map satisfying the above is called unital.
In Ref. [15, 24], depolarization noise is considered, $D_{p}(\cdot)=(1-p)(\cdot)+p\frac{\mathds{1}}{2}$,
$\widetilde{M}_{i}=(1-\frac{p}{2})|i\rangle\langle i|+\frac{p}{2}|i\oplus 1\rangle\langle i\oplus 1|$ (3)
where $\oplus$ is a bitwise addition. We have used $\mathds{1}=|0\rangle\langle 0|+|1\rangle\langle 1|$. A depolarization noise may be regarded as a mixture of all types of noise. In Ref. [24], it has been shown that measurements with depolarization noise cannot be used to distill entanglement from weakly entangled states, see also Fig. 3.
@@
The purification protocol works as follows. We start with general noise in Eq. (2) and show that depolarization noise corresponds to the worst-case for entanglement distillation. A high-dimensional extension is shown in Supplemental Material. Let $\rho$ denote a qubit state on which a measurement is performed. A noiseless measurement shows probabilities $\mathrm{tr}[\rho M_{i}]$, and a noisy one finds $\mathrm{tr}[\rho\widetilde{M}_{i}]$ for $i=0,1$. We here devise a protocol to obtain error-free statistics $\mathrm{tr}[\rho M_{i}]$ with noisy measurements and additional qubits as resources.
@@
Figure 1: A measurement is noisy (red), see Eq. (3). An error-suppressing detector (blue) is devised with $n-1$ additional qubits and a collective CNOT $V_{n}$ in Eq. (4).
@@
Measurements on $n$ qubits are noisy. Suppose that $n$-bit outcomes are obtained, $x^{n}=(x_{1},x_{2},\cdots,x_{n})$ where $x_{k}\in\{0,1\}$ for $k=1,\cdots,n$. A POVM element describing the outcomes correspond to $\widetilde{M}_{x^{n}}=\widetilde{M}_{x_{1}}\otimes\cdots\otimes\widetilde{M}_{x_{n}}$. We have the probability of outcome $x^{n}$,
@@
We accept measurement results only when all outcomes are identical, i.e., either $0^{n}$ or $1^{n}$; otherwise, discard. This approach has its origin in the advantage distillation [26, 27, 28, 7, 30, 31]. Note also that a similar approach can tolerantly deal with losses in the preparation of cluster states [32]. Once outcomes are accepted, a probability of outcome $i\in\{0,1\}$,
@@
which converges to a probability obtained by a noiseless measurement.
To see that a probability in Eq. (7) is arbitrarily close to a result of a noiseless measurement, let us rewrite Eq. (6) in terms of an effective POVM element for an outcome $x_{1}\in\{0,1\}$,
@@
$\mathrm{where}$ $\widetilde{Q}_{x_{1}}^{(n)}=\langle 0_{2}\cdots 0_{n}|V_{n}^{\dagger}\widetilde{M}_{x^{n}}V_{n}|0_{2}\cdots 0_{n}\rangle.$
To see the derivation of Eq. (8) from Eq. (6), we have used $\mathrm{tr}[AB]=\mathrm{tr}[BA]$ for Hermitian operators $A$ and $B$ so that
$\mathrm{tr}[V_{n}~{}\rho\otimes|0\rangle\langle 0|^{\otimes n-1}~{}V_{n}^{\dagger}\widetilde{M}_{x^{n}}]$ $=$ $\mathrm{tr}[~{}\rho\otimes|0\rangle\langle 0|^{\otimes n-1}~{}V_{n}^{\dagger}\widetilde{M}_{x^{n}}V_{n}]$ (9)
$=$ $\mathrm{tr}[~{}\rho\langle 0_{2}\cdots 0_{n}|V_{n}^{\dagger}\widetilde{M}_{x^{n}}V_{n}|0_{2}\cdots 0_{n}\rangle].$
It holds that $V_{n}|0_{2}\cdots 0_{n}\rangle=|0^{n}\rangle\langle 0|+|1^{n}\rangle\langle 1|$, from which corresponding POVM elements when the outcomes are identical, $x^{n}\in\{0^{n},1^{n}\}$, can be found as follows.
Suppose that as the consequence of unwanted interaction in Eq. (2), a noisy measurement is given by,
$\widetilde{M}_{0}=\left[{\begin{array}[]{cc}1-s&t\\ t&s\\ \end{array}}\right]~{}~{}\mathrm{and}~{}~{}\widetilde{M}_{1}=\left[{\begin{array}[]{cc}s&-t\\ -t&1-s\\ \end{array}}\right]$ (14)
where $s\in[0,1/2)$ and $t\in[-\alpha,\alpha]$ with $\alpha=\sqrt{s(1-s)}$ since $\widetilde{M}_{i}\geq 0$. Let us consider cases when $n$ outcomes are identical, i.e., $x_{1}=\cdots=x_{n}\in\{0,1\}$. Let $i\in\{0,1\}$ denote the value accepted in the measurement. We may characterize POVM elements for an outcome $x_{1}=i\in\{0,1\}$ in Eq. (6) are given by
$\widetilde{Q}_{i}^{(n)}$ $=$ $r_{0}^{(n)}M_{i}+r_{1}^{(n)}M_{i\oplus 1}+((-1)^{i}t)^{n}(|0\rangle\langle 1|+|1\rangle\langle 0|)~{}~{}$
$\mathrm{where}$ $r_{0}^{(n)}=(1-s)^{n}~{}\mathrm{and}~{}r_{1}^{(n)}=s^{n}.$ (15)
@@
The derivations are detailed in Supplemental Material. The probability in Eq. (7) can be rewritten by defining POVM elements $Q_{i}^{(n)}$ on the first qubit when all outcomes are identical:
$q^{(n)}(i)$ $=$ $\mathrm{tr}[\rho Q_{i}^{(n)}]~{}\mathrm{where}~{}Q_{i}^{(n)}=\frac{\widetilde{Q}_{i}^{(n)}}{\sum_{i=0,1}p(i^{n})}~{}~{}~{}~{}~{}$ (16)
One can find $n\geq 1$ such that $Q_{i}^{(n)}$ is $\epsilon$-close to a clean one $M_{i}$;
$\frac{1}{2}\|Q_{i}^{(n)}-M_{i}\|_{1}\leq\epsilon,$
where $\|\cdot\|_{1}$ denotes a trace norm. i.e., $\|X\|_{1}=\mathrm{tr}\sqrt{X^{\dagger}X}$. We stress that for realistic cases, in particular, for depolarizing noise, where we have an error rate less than $10\%$, i.e., $s\leq 0.1$, one additional qubit, i.e., $n=2$, suffices to get $\epsilon=2\times 10^{-2}$ for which the probability of having identical outcomes is over $0.8$. The analysis is detailed in Supplemental Material.
III Distilling Entanglement
We now consider the distillation of entanglement with noisy local operations. We consider the distillation protocol in Ref. [11] because of the following reasons. First, it can distill ebits from all two-qubit entangled states. Second, a single CNOT gate by each party suffices. Generalizations of the protocol can be considered, e.g., [33, 34], with a greater number of CNOT gates. However, an improvement is not immediate. The generalizations show enhancement in a limited range of an initial fidelity $F$ in Eq. (19). In addition, a greater number of CNOT gates are experimentally demanding. The protocol in Ref. [11] is the most robust and experimentally feasible, e.g., [17, 18, 35, 22].
@@ III Distilling Entanglement
Let us summarize the distillation protocol [11]. It begins with copies of unknown bipartite states $\rho^{\otimes N}$ for a large $N$. Two parties apply state twirling by LOCC, which transforms a shared state to an isotropic state at a single-copy level as follows,
$\rho\mapsto\rho_{F}=\sum_{k=1}^{m}p_{k}U_{k}\otimes U_{k}^{\ast}\rho(U_{k}\otimes U_{k}^{\ast})^{\dagger}$
where denotes the complex conjugation with respect to the computational basis. With unitary transformations that form a quantum $2$-design $\{p_{K},U_{k}\}_{k=1}^{m}$ [36, 37]. An isotropic state can be characterized by a single parameter $F$ as follows,
@@
where $F=\langle\phi^{+}|\rho_{F}|\phi^{+}\rangle$ is called a singlet fraction.
In what follows, the goal is to distill an ebit in Eq. (1) from copies of isotropic states. In a realistic scenario, one should admit some fraction of errors in achieving an ebit, e.g., one may refer to a scheme for robust self-testing for an ebit, e.g., [38]. For clarification, let us introduce an $\epsilon$-ball of an ebit
$B_{\epsilon}(|\phi^{+}\rangle)=\{\sigma\in S(\mathcal{H}\otimes\mathcal{H})~{}:~{}\frac{1}{2}\|\sigma-|\phi^{+}\rangle\langle\phi^{+}|\|_{1}\leq\epsilon\}$ (18)
where $S(\mathcal{H})$ denotes the set of states on a Hilbert space $\mathcal{H}$. Then, one can distinguish a state $\sigma\in B_{\epsilon}(|\phi^{+}\rangle)$ and an ideal instance $|\phi^{+}\rangle$ with a probability at most $1/2+\epsilon$, which is $\epsilon$ better than the random. From the well-known relation between the state fidelity and the trace distance [39], we have for a state $\sigma\in B_{\epsilon}(|\phi^{+}\rangle)$,
$F(\sigma,|\phi^{+}\rangle)$ $=$ $\langle\phi^{+}|\sigma|\phi^{+}\rangle$
$=$ $1-\frac{1}{2}\|\sigma-|\phi^{+}\rangle\langle\phi^{+}|\|_{1}$
which is at least $1-\epsilon$. Throughout, distilling an ebit signifies a state belonging to $B_{\epsilon}(|\phi^{+}\rangle)$ admitting a precision up to $\epsilon$ so that we may achieve an ebit with a fidelity higher than $1-\epsilon$.
III.1 Noiseless scenario
Let us follow the steps for copies of isotropic states, see also Fig. 2. i) Two parties take two copies $\rho_{F}^{A_{1}B_{1}}\otimes\rho_{F}^{A_{2}B_{2}}$, and apply bilateral CNOT gates, $V_{2}^{A_{1}A_{2}}$ and $V_{2}^{B_{1}B_{2}}$, see Eq. (4). ii) Two parties perform measurements in the second register $A_{2}B_{2}$ in the computational basis $\{M_{0},M_{1}\}$. iii) If outcomes are equal, a state remaining in the first register $A_{1}B_{1}$ is accepted. Otherwise, they abort the protocol and start step i) with isotropic states. Once states are accepted, two parties resume the protocol with accepted copies from the state twirling.
Suppose that measurement outcomes in the second register are identical and the copy in the first register is accepted. The singlet fidelity of the accepted state is given by
@@ III.1 Noiseless scenario
Figure 2: The protocol for distilling entanglement works with a pair of isotropic states $\rho_{F}^{\otimes 2}$ distributed to Alice and Bob, see Eq. (17). Then, each party applies a CNOT gate and performs a measurement on a qubit in the second register. A resulting state in the first register is accepted only when the measurement outcomes are identical. Then, an ebit can be distilled whenever $F>1/2$. When measurements in the second register are noisy, an ebit can be distilled for a limited range $F>L$ for some lower bound $L>1/2$. Then, two parties apply error-suppressing detectors and recover a lower bound $L$ sufficiently close to $1/2$. For an error rate up to $5\%$, one additional qubit on each party immediately allows entanglement distillation for an initial singlet fraction up to $F>0.505$.
III.2 Noisy scenario
In Ref. [24], noisy LOCC are considered in the distillation protocol, particularly, measurements in the second register $A_{2}B_{2}$ suffering from a depolarization noise, see Eq. (3). It turns out that the protocol with noisy LOCC cannot distill ebits from all two-qubit entangled states. For instance, for a depolarization noise with $p=0.1$ in Eq. (3), meaning an error rate is $5\%$, entanglement can be distilled only when $F>0.617$ [24]. No entanglement can be distilled if $p>1-1/\sqrt{2}\approx 0.29$: noise is crucial in the distillation of entanglement.
In what follows, we incorporate the protocol for purifying noisy measurements to the distillation of entanglement, see Fig. 2. The goal is primarily to improve a lower bound of an initial fidelity, denoted by $F>L$, such that a resulting fidelity increases $F^{{}^{\prime}}>F$. That is, $L=1/2$ for noiseless LOCC and $L>1/2$ for noisy LOCC with a depolarization noise on measurements [24], and we aim to maintain $L=1/2$ even if LOCC are noisy. As a byproduct, the improvement also allows us to optimize the resources for entanglement distillation.
For generality, we do not restrict to measurements with a depolarization noise in Eq. (3), considered in Ref. [24]. We begin with measurements containing general noise in Eq. (14) and seek a lower bound $L^{(n,m)}$ with $n-1$ and $m-1$ target qubits of Alice and Bob, respectively, for purification of noisy measurements.
We now remark that noisy measurements with a depolarization noise in Eq. (3), that is, particular instances with $t=0$ in Eq. (14), lead to the worst-case scenario. To be precise, let us suppose that an error rate $s$ is given in a noisy measurement in Eq. (14) and apply them to the distillation of entanglement. A lower bound for the distillability may depend on parameters $s$ and $t$, denoted by $L^{(1,1)}(s,t)$ without the purification protocol yet, turns out to be the highest for $t=0$, which corresponds to a depolarization noise in Eq. (3), see Fig. 4. The proof is detailed in Supplemental Material. Hence, for a given error rate $s$, we have,
$\max_{s}L^{(1,1)}(s,t)=L^{(1,1)}(s,0),$ (20)
The results in Ref. [24] consider the worst-case scenario.
Figure 3: The robustness of the purification protocol is demonstrated. When a measurement is noiseless, an ebit can be distilled from entangled states having $F>1/2$. (a) For a measurement error $p=0.1$, the condition for distilling entanglement is more stringent $F>0.617$, which is the case $n=m=1$; see the red dot in the plot. With an additional qubit, $n=m=2$, the condition improves up to $F>0.505$; see the yellow dot. Thus, an ebit can be distilled from a larger set of entangled states. (b) For a larger measurement error $p=0.2$, the condition is even worse, $F>0.779$; see the red dot. With an additional qubit, an ebit can be distilled for $F>0.53$; see the yellow dot. Then, as it is discussed in Eq. (26), measurements may have different error rates, for which let $p_{i}$ denotes an error rate of the $i$-th measurement. The robustness is demonstrated by varying a parameter $\delta_{i}$ in the range $p_{i}\in[p-\delta_{i},p+\delta_{i}]$ for $\delta_{i}\in[0,0.075]$. With randomly chosen sequences $\delta_{i}$, the distillation protocol works faithfully to distill ebits; we repeat numerical demonstration and show the range, see the grey area. The robustness also holds for other values of $p$.